# On Erlang mixture approximations for differential equations

### v2

@@ 5.2 Convergence of the roots
for $\operatorname{Re}\lambda>-\bar{\rho}$. Furthermore, for any $\lambda$ for which $\operatorname{Re}\lambda>-\rho$, we can choose $\bar{a}$ large enough that $\operatorname{Re}\lambda>-\bar{\rho}>-\rho$ whenever $a_{i}>\bar{a}$ for $i=1,\ldots,n_{z}$. Consequently, the derivative exists when $\operatorname{Re}\lambda>-\rho$. Finally, as the determinant of a holomorphic function is also holomorphic and the remaining terms in the argument of the determinant in the expression for $\hat{P}$ in (83) are constant and linear in $\lambda$, $\hat{P}$ is holomorphic when $\operatorname{Re}\lambda>\rho$.
Theorem 5.14.
Let each element of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ be a regular and exponentially bounded kernel such that there exist $L,\rho\in\mathbb{R}_{>0}$ for which $\alpha_{i}(t)\leq Le^{-\rho t}$ for all $t\in\mathbb{R}_{\geq 0}$ and $i=1,\ldots,n_{z}$, and let element $i$ of $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ be an Erlang mixture approximation of infinite order with rate parameter $a_{i}\in\mathbb{R}_{>0}$ and coefficients defined by (17) based on $\alpha_{i}$. Then, for every isolated root $\lambda\in\mathbb{C}$ of the characteristic function $P:\mathbb{C}\rightarrow\mathbb{C}$ for which $\operatorname{Re}\lambda\geq\lambda_{\min}$ where $\rho<\lambda_{\min}\in\mathbb{R}_{\leq 0}$, there are roots of the approximate characteristic function $\hat{P}$$\mathbb{C}\rightarrow\mathbb{C}$ defined in (83) that converge to $\lambda$.
Proof 5.15.
According to Lemma 5.12 and 5.10, respectively, $\hat{P}$ is holomorphic and converges to $P$ for $\lambda\in\mathbb{C}$ with real part larger than or equal to $\lambda_{\min}$. Consequently, the convergence of the roots follows from Hurwitz’ convergence theorem [19, Thm. 5.6.4] since the roots are isolated by assumption.
@@ 5.3 Error dynamics
In this section, we analyze the difference between the solution to the approximate system of ODEs (73) and that of the original system of DDEs (2b)–(3), and we show that for a given time $t\in[t_{0},t_{f}]$, it converges to zero as the Erlang mixture approximation becomes more accurate.
@@
where $\hat{x}$ and $\hat{z}$ are the solution to the approximate system of ODEs (73) and $x$ and $z$ are the solution to the original system of DDEs in (2b)–(3). Both $\hat{x}$ and $x$ are assumed to satisfy the initial condition (2a). The state error satisfies the system of DDEs
@@
Lemma 5.16.
The elements of the memory state error, $e_{z}:\mathbb{R}\rightarrow\mathbb{R}^{n_{z}}$ are bounded by
(108) $|e_{z,i}(t)|$ $\leq\int_{-\infty}^{t}\hat{\alpha}_{i}(t-s)|e_{r,i}(s)|\,\mathrm{d}s+K_{r}\int_{-\infty}^{t}|e_{\alpha,i}(s)|\,\mathrm{d}s\,r_{i}(s),$ $i$ $=1,\ldots,n_{z},$
for $t\in[t_{0},t_{f}]$.
Proof 5.17.
First, we rewrite the error of the $i$’th memory state by adding and subtracting the integral of $\hat{\alpha}_{i}(t-s)r_{i}(s)$ (i.e., a mixed term):
(109) $e_{z,i}(t)$ $=\int_{-\infty}^{t}\hat{\alpha}_{i}(t-s)\hat{r}_{i}(s)\,\mathrm{d}s-\int_{-\infty}^{t}\alpha_{i}(t-s)r_{i}(s)\,\mathrm{d}s$
$=\int_{-\infty}^{t}\hat{\alpha}_{i}(t-s)e_{r,i}(s)\,\mathrm{d}s+\int_{-\infty}^{t}e_{\alpha,i}(t-s)r_{i}(s)\,\mathrm{d}s.$
Next, we bound the absolute value of the error using a triangle inequality and the boundedness of the delayed variable, $r$:
(110) $|e_{z,i}(t)|$ $\leq\int_{-\infty}^{t}\hat{\alpha}_{i}(t-s)|e_{r,i}(s)|\,\mathrm{d}s+\int_{-\infty}^{t}|e_{\alpha,i}(t-s)||r_{i}(s)|\,\mathrm{d}s$
$\leq\int_{-\infty}^{t}\hat{\alpha}_{i}(t-s)|e_{r,i}(s)|\,\mathrm{d}s+\int_{-\infty}^{t}|e_{\alpha,i}(t-s)|\,\mathrm{d}s\sup_{s\in(-\infty,t]}r_{i}(s).$
By assumption (see Section 3), $K_{r}=\sup_{t\in(-\infty,t_{f}]}r_{i}(t)$.
Corollary 5.18.
The initial memory state error is bounded by
(111) $|e_{z,i}(t_{0})|$ $\leq K_{r}\int_{-\infty}^{t_{0}}|e_{\alpha,i}(t-s)|\,\mathrm{d}s.$
Proof 5.19.
For $t\leq t_{0}$, $\hat{x}(t)=x(t)$ such that $e_{r}(t)=0$.
Theorem 5.20.
If the original system of DDEs (2b)–(3) and the approximate system (71) are both locally asymptotically stable around a steady state, $\bar{x}$, (see Corollary 3.5 and Theorem 5.4), so is the system (107b) describing the state error. Otherwise, (107b) is not asymptotically stable.
Proof 5.21.
The steady states of the original and approximate system are identical (see Theorem 5.2). The linearized system corresponding to the DDE (107b) describing the state error is (see also the identity (109))
(112) $\dot{e}_{x}(t)$ $=Fe_{x}(t)+G\left(\int_{-\infty}^{t}\hat{\alpha}(t-s)\odot He_{x}(s)\,\mathrm{d}s+\int_{-\infty}^{t}e_{\alpha}(t-s)\odot HX(s)\,\mathrm{d}s\right).$
@@
(113) $\begin{bmatrix}\dot{X}(t)\\ \dot{e}_{x}(t)\end{bmatrix}=\begin{bmatrix}F\\ &F\end{bmatrix}\begin{bmatrix}X(t)\\ e_{x}(t)\end{bmatrix}\\ +\begin{bmatrix}G\\ &G\end{bmatrix}\int_{-\infty}^{t}\begin{bmatrix}\operatorname{diag}\alpha(t-s)\\ \operatorname{diag}e_{\alpha}(t-s)&\operatorname{diag}\hat{\alpha}(t-s)\end{bmatrix}\begin{bmatrix}H\\ &H\end{bmatrix}\begin{bmatrix}X(s)\\ e_{x}(s)\end{bmatrix}\,\mathrm{d}s.$
@@
$=\det(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\alpha(s)\,\mathrm{d}sH)\det(F-\lambda I+G\int_{0}^{\infty}e^{-\lambda s}\operatorname{diag}\hat{\alpha}(s)\,\mathrm{d}sH)$
@@
where $P$ and $\hat{P}$ are the characteristic functions corresponding to the original system of DDEs (2b)–(3) and the approximate system (71), respectively. Consequently, the roots of $P$ and $\hat{P}$ are also roots of the above characteristic function, and if their real parts are negative, the augmented system is asymptotically stable and the error dynamics are locally asymptotically stable.
Theorem 5.22.
Let the elements of $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ be exponentially bounded regular kernels, and let the $i$’th element of $\hat{\alpha}:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}^{n_{z}}$ be the infinite-order Erlang mixture kernel obtained with the rate parameter $a_{i}\in\mathbb{R}_{>0}$ and the coefficients (17) based on $\alpha_{i}$. Then, for given $t\in[t_{0},t_{f}]$ and for every $\epsilon\in\mathbb{R}_{>0}$, there exists an $\bar{a}\in\mathbb{R}_{>0}$ such that
(115) $\|\hat{x}(t)-x(t)\|_{\infty}$ $\leq\epsilon w(t)\leq\epsilon w(t_{f}),$
when $a_{i}>\bar{a}$ for $i=1,\ldots,n_{z}$. The auxiliary function $w:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is the solution to the DDE
@@
where $L_{x}\in\mathbb{R}_{>0}$ and $L_{z}\in\mathbb{R}_{>0}$ are the Lipschitz constants of the first and second arguments of $f$, and $L_{r}\in\mathbb{R}_{>0}$ is the Lipschitz constant of $h$. Furthermore, $K_{r}$ is the supremum of $r$ in the time interval $(-\infty,t_{f}]$ (see also the assumptions in Section 3).
Proof 5.23.
According to Theorem 4.37, we can choose $\bar{a}$ large enough that
(117) $\int_{0}^{\infty}\|e_{\alpha}(t)\|_{\infty}\,\mathrm{d}t$ $<\epsilon$
when $a_{i}>\bar{a}$ for $i=1,\ldots,n_{z}$. Furthermore, since 1) the kernel $\alpha$ is non-negative and its integral is one, 2) the delayed variable, $r$, is assumed to be bounded, and 3) the functions $f$ and $h$ are assumed to be differentiable (and thus also Lipschitz continuous), the result follows from a proof similar to that of Theorem 1 in the paper by Guglielmi and Hairer [18]. For completeness, we present the details of the proof in Section E of the supplementary material. Finally, the second inequality in (115) follows from the non-negativity of the right-hand side of the DDE (116b).
@@ 6 Algorithm
In Section 4 and 5, we have shown that the DDEs (2b)–(3) can be approximated arbitrarily well by the ODEs (73). However, the convergence results presume that the order, $M$, of the Erlang mixture approximation is infinite. In that case, the ODEs are infinite-dimensional. Therefore, and because the infinite number of coefficients, $\{c_{m}\}_{m=0}^{\infty}$, must sum to one, we truncate the series in the approximation to a finite number of terms, $M+1$.
Furthermore, we describe an algorithm for determining the coefficients and the rate parameter in an $M$’th order Erlang mixture approximation of a regular kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}_{\geq 0}$, for given $M\in\mathbb{N}_{\geq 0}$. For simplicity, we only consider scalar kernels in this section. The algorithm consists of two steps: First, we identify a domain in which we approximate the kernel and then, we use a least-squares approach to determine the coefficients and the rate parameter. Furthermore, we describe a reference approach based on the theoretical expressions (17) for the coefficients. Finally, we describe the implementation of the algorithm and the simulation of the original DDEs (2b)–(3b) and the approximate system of ODEs (73).
@@ 6.1 Domain
In order to determine the approximation domain, we introduce $\beta:\mathbb{R}_{\geq 0}\rightarrow[0,1]$ given by
(118) $\beta(t)$ $=\int_{0}^{t}\alpha(s)\,\mathrm{d}s,$
which is analogous to the cumulative distribution function of a random variable with probability density function $\alpha$. As $\alpha$ is non-negative, $\beta$ is non-decreasing, and for a given threshold, $\epsilon\in\mathbb{R}_{>0}$, we choose $t_{h}\in\mathbb{R}_{\geq 0}$ as the solution to the equation $1-\beta(t)=\epsilon$ and approximate $\alpha$ in the interval $0,t_{h}]$. We use bisection to determine $t_{h}$ approximately. We assume that an initial interval, $t_{0}^{\ell},t_{0}^{u}]$, is available and that it contains the root, i.e., that $1-\beta(t_{0}^{\ell})>\epsilon$ and $1-\beta(t_{0}^{u})<\epsilon$. If that is not the case, the interval can be shifted or scaled such that it does. Next, the $k$’th iteration has converged and an approximate solution has been found if $|1-\beta(t_{k}^{m})-\epsilon|$ is below a specified threshold, where $t_{k}^{m}=\frac{1}{2}(t_{k}^{\ell}+t_{k}^{u})$ is the midpoint. Otherwise, the next iteration considers the lower half of the interval, $t_{k+1}^{\ell},t_{k+1}^{u}]=[t_{k}^{\ell},t_{k}^{m}]$, if $1-\beta(t_{k}^{m})<\epsilon$. Similarly, if $1-\beta(t_{k}^{m})>\epsilon$, the next iteration considers the upper half of the interval, $t_{k+1}^{\ell},t_{k+1}^{u}]=[t_{k}^{m},t_{k}^{u}]$.
@@
If the function $\beta$ given by (118) cannot be derived analytically for a given kernel, $\alpha$, it can be approximated using numerical quadrature (e.g., with Matlab’s function integral).
@@
The equation $1-\beta(t)=\epsilon$ can also be solved using a gradient-based approach, e.g., Newton’s method. However, $1-\beta(t)$ rounds to zero in finite-precision arithmetic when $t$ is too large, e.g., in a computer implementation. Consequently, Newton iterations may stall if an iterate happens to become too large.

### v3

@@ 4.3 Convergence of the roots
for $\operatorname{Re}\lambda>-\bar{\rho}$. Furthermore, for any $\lambda$ for which $\operatorname{Re}\lambda>-\rho$, we can choose $\bar{a}$ and $\bar{M}$ (for which $\bar{M}/\bar{a}\rightarrow\infty$ as $\bar{a}\rightarrow\infty$) large enough that $\operatorname{Re}\lambda>-\bar{\rho}>-\rho$ whenever $a_{i}>\bar{a}$ and $M_{i}>\bar{M}$ for $i=1,\ldots,n_{z}$. Consequently, the derivative exists when $\operatorname{Re}\lambda>-\rho$. Finally, as the determinant of a holomorphic function is also holomorphic and the remaining terms in the argument of the determinant in the expression for $\hat{P}$ in (38) are constant and linear in $\lambda$, $\hat{P}$ is holomorphic when $\operatorname{Re}\lambda>\rho$.
Proof 4.22 (Proof of Theorem 4.4(b)).
According to Lemma 4.20 and 4.18, respectively, $\hat{P}$ is holomorphic and converges uniformly to $P$ for $\lambda\in\mathbb{C}$ with real part larger than or equal to $\lambda_{\min}$. Consequently, the convergence of the roots of $\hat{P}$ follows from Hurwitz’ convergence theorem [19, Thm. 5.6.4] since the roots are isolated by assumption.
@@ 4.4 Error dynamics
In this section, we analyze the difference between the solution to the IVP (19)–(20) involving approximate ODEs and that of the original IVP (3)–(4) involving DDEs, and we show that for a given time $t\in[t_{0},t_{f}]$, it converges to zero as the Erlang mixture approximation becomes more accurate.
@@
where $\hat{x}$ and $\hat{z}$ are the solution to the IVP (19)–(20) involving ODEs and $x$ and $z$ are the solution to the original IVP (3)–(4) involving DDEs. The state error satisfies the IVP
@@
Proof 4.23 (Proof of Theorem 4.4(c)).
The steady states of the original and approximate system are not identical (see Lemma 4.11). Consequently, the steady state of the error, $\bar{e}_{x}\in\mathbb{R}^{n_{x}}$, will not be zero. The linearized system corresponding to the DDE (60b) describing the state error is
(61) $\dot{E}_{x}(t)$ $=\hat{F}\hat{X}(t)+\hat{G}\int_{-\infty}^{t}\hat{\alpha}(t-s)\odot\hat{H}\hat{X}(s)\,\mathrm{d}s-\left(FX(t)+G\int_{-\infty}^{t}\alpha(t-s)\odot HX(s)\,\mathrm{d}s\right),$
where $E_{x}:\mathbb{R}\rightarrow\mathbb{R}^{n_{x}}$ is the deviation from the steady state error, i.e., $E_{x}(t)=e_{x}(t)-\bar{e}_{x}$, and $\hat{X},X:\mathbb{R}\rightarrow\mathbb{R}^{n_{x}}$ are the deviations of the state variables from their respective steady states. Next, we rewrite the system by adding and subtracting terms such that $\hat{X}$ does not appear on the right-hand side:
(62) $\dot{E}_{x}(t)$ $=\hat{F}E_{x}(t)+e_{F}X(t)+G\int_{-\infty}^{t}\hat{\alpha}(t-s)\odot\hat{H}E_{x}(s)\,\mathrm{d}s+\hat{G}\int_{-\infty}^{t}\hat{\alpha}(t-s)\odot e_{H}X(s)\,\mathrm{d}s$
$+\hat{G}\int_{-\infty}^{t}e_{\alpha}(t-s)\odot HX(s)\,\mathrm{d}s+e_{G}\int_{-\infty}^{t}\alpha(t-s)\odot HX(s)\,\mathrm{d}s.$
@@
(63) $\begin{bmatrix}\dot{X}(t)\\ \dot{E}_{x}(t)\end{bmatrix}=\begin{bmatrix}F\\ e_{F}&\hat{F}\end{bmatrix}\begin{bmatrix}X(t)\\ E_{x}(t)\end{bmatrix}\\ +\begin{bmatrix}G\\ e_{G}&\hat{G}\end{bmatrix}\int_{-\infty}^{t}\begin{bmatrix}\operatorname{diag}\alpha(t-s)\\ \operatorname{diag}e_{\alpha}(t-s)&\operatorname{diag}\hat{\alpha}(t-s)\end{bmatrix}\begin{bmatrix}H\\ e_{H}&\hat{H}\end{bmatrix}\begin{bmatrix}X(s)\\ E_{x}(s)\end{bmatrix}\,\mathrm{d}s.$
@@
[…]
@@
where $P$ and $\hat{P}$ are the characteristic functions corresponding to the original system of DDEs (3b)–(4) and the approximate system of ODEs (19), respectively. Consequently, the roots of $P$ and $\hat{P}$ are also roots of the above characteristic function, and if their real parts are negative, the augmented system is asymptotically stable and the error dynamics are locally asymptotically stable.
4.5 Uniform convergence
In this section, we prove Theorem 4.4(d) on the uniform convergence of the solution to the IVP (19)–(20) involving ODEs to the solution of the original IVP (3)–(4) involving DDEs.
Proof 4.24 (Proof of Theorem 4.4(d)).
First, we prove that the solution, $\hat{x}$, to the IVP (19)–(20) with ODEs converges uniformly to that of the IVP (3)–(4) with DDEs. According to Proposition 4.2(d), we can choose $\bar{a}$ and $\bar{M}$ large enough that
(65) $\int_{0}^{\infty}\|e_{\alpha}(t)\|_{\infty}\,\mathrm{d}t$ $<\epsilon$
when $a_{i}>\bar{a}$ and $M_{i}>\bar{M}$ for $i=1,\ldots,n_{z}$ provided that $\bar{M}/\bar{a}\rightarrow\infty$ as $\bar{a}\rightarrow\infty$. Furthermore, since 1) the delayed variable, $r$, is assumed to be bounded and 2) the functions $f$ and $h$ are assumed to be differentiable (and thus also Lipschitz continuous), the result follows from a proof similar to that of Theorem 1 in the paper by Guglielmi and Hairer [18]. For completeness, we present the details of the proof in Section SM3 of the supplementary material. Specifically,
(66) $\|e_{x}(t)\|_{\infty}$ $\leq\epsilon w(t)\leq\epsilon w(t_{f}),$
where the auxiliary function $w:\mathbb{R}\rightarrow\mathbb{R}_{\geq 0}$ is the solution to the DDE
@@ 4.5 Uniform convergence
and $L_{x}\in\mathbb{R}_{>0}$ and $L_{z}\in\mathbb{R}_{>0}$ are the Lipschitz constants of the first and second arguments of $f$, and $L_{r}\in\mathbb{R}_{>0}$ is the Lipschitz constant of $h$. Furthermore, $K_{r}$ is the supremum of the absolute value of $r$ in the time interval $(-\infty,t_{f}]$ (see also the assumptions in Section 3). The second inequality in (66) follows from the non-negativity of the right-hand side of the DDE (67b) and ensures that the convergence is uniform on the time interval $t_{0},t_{f}]$. In conclusion, the state error, $e_{x}$, can be made arbitrarily small by making the kernel error sufficiently small.
Next, we prove that $\hat{z}$ and $\hat{r}$ converge uniformly. As $h$ is differentiable, $|e_{r,i}(t)|\leq L_{r}|e_{x,i}(t)|\leq L_{r}\epsilon w(t_{f})$, and therefore, the convergence of $\hat{r}$ is uniform. We bound the integral of the absolute value of the $i$’th Erlang mixture approximation by
(68) $\int_{0}^{\infty}|\hat{\alpha}_{i}(t)|\,\mathrm{d}t$ $\leq\sum_{m=0}^{M}|c_{m}|\int_{0}^{\infty}\ell_{m}(t)\,\mathrm{d}t=\sum_{m=0}^{M}|c_{m}|\leq\sum_{m=0}^{M}\int_{s_{m}}^{s_{m+1}}|\alpha(t)|\,\mathrm{d}t$
$=\int_{0}^{s_{M+1}}|\alpha(t)|\,\mathrm{d}t\leq\rho(\infty),$
which is finite by assumption. Using these two bounds, we derive a bound on the absolute memory state error:
(69) $|e_{z,i}(t)|$ $=\left|\int_{-\infty}^{t}\hat{\alpha}_{i}(t-s)\hat{r}_{i}(s)\,\mathrm{d}s-\int_{-\infty}^{t}\alpha_{i}(t-s)r_{i}(s)\,\mathrm{d}s\right|$
$=\left|\int_{-\infty}^{t}\hat{\alpha}_{i}(t-s)e_{r,i}(s)\,\mathrm{d}s+\int_{-\infty}^{t}e_{\alpha,i}(t-s)r_{i}(s)\,\mathrm{d}s\right|$
$\leq\int_{-\infty}^{t}|\hat{\alpha}_{i}(t-s)||e_{r,i}(s)|\,\mathrm{d}s+\int_{-\infty}^{t}|e_{\alpha,i}(t-s)||r_{i}(s)|\,\mathrm{d}s$
$\leq\int_{-\infty}^{t}|\hat{\alpha}_{i}(t-s)|\,\mathrm{d}s\sup_{s\in(-\infty,t]}|e_{r,i}(s)|+\int_{-\infty}^{t}|e_{\alpha,i}(t-s)|\,\mathrm{d}s\sup_{s\in(-\infty,t]}|r_{i}(s)|$
$\leq(\rho(\infty)L_{r}w(t_{f})+K_{r})\epsilon.$
From the first to the second equality, we add and subtract the integral of $\hat{\alpha}_{i}(t-s)r_{i}(s)$ (i.e., a mixed term). In conclusion, we can bound the pointwise error, and the bound is independent of $t$ for $t\in(-\infty,t_{f}]$, i.e., the convergence is uniform over that interval.
@@ 5 Algorithm
In this section, we describe an algorithm for determining the coefficients and the rate parameter in an $M$’th order Erlang mixture approximation of a regular kernel, $\alpha:\mathbb{R}_{\geq 0}\rightarrow\mathbb{R}$, for given $M\in\mathbb{N}_{\geq 0}$. For simplicity, we only consider scalar kernels in this section. The algorithm consists of two steps: First, we identify a domain in which we approximate the kernel and then, we use a least-squares approach to determine the coefficients. Furthermore, we describe a reference approach based on the theoretical expressions in (15) for the coefficients. Finally, we describe the implementation of the algorithm and the simulation of the original DDEs (3b)–(4) and the approximate system of ODEs (19).
@@ 5.1 Domain
In order to determine the approximation domain, we choose $t_{h}\in\mathbb{R}_{\geq 0}$ as the solution to the equation $\rho(\infty)-\rho(t)=\epsilon$ for a given threshold, $\epsilon\in\mathbb{R}_{>0}$, and approximate $\alpha$ in the interval $0,t_{h}]$. We use bisection to determine $t_{h}$ approximately. We assume that an initial interval, $t_{0}^{\ell},t_{0}^{u}]$, is available and that it contains the root, i.e., that $\rho(\infty)-\rho(t_{0}^{\ell})>\epsilon$ and $\rho(\infty)-\rho(t_{0}^{u})<\epsilon$. If that is not the case, the interval can be shifted or scaled such that it does. Next, the $k$’th iteration has converged and an approximate solution has been found if $|\rho(\infty)-\rho(t_{k}^{m})-\epsilon|$ is below a specified threshold, where $t_{k}^{m}=\frac{1}{2}(t_{k}^{\ell}+t_{k}^{u})$ is the midpoint. Otherwise, the next iteration considers the lower half of the interval, $t_{k+1}^{\ell},t_{k+1}^{u}]=[t_{k}^{\ell},t_{k}^{m}]$, if $\rho(\infty)-\rho(t_{k}^{m})<\epsilon$. Conversely, if $\rho(\infty)-\rho(t_{k}^{m})>\epsilon$, the next iteration considers the upper half of the interval, $t_{k+1}^{\ell},t_{k+1}^{u}]=[t_{k}^{m},t_{k}^{u}]$.
@@
If the function $\rho$ in (5) cannot be derived analytically for a given kernel, $\alpha$, it can be approximated using numerical quadrature (e.g., with Matlab’s function integral).
@@
The equation $\rho(\infty)-\rho(t)=\epsilon$ can also be solved using a gradient-based approach, e.g., Newton’s method. However, $\rho(\infty)-\rho(t)$ rounds to zero in finite-precision arithmetic when $t$ is too large, e.g., in a computer implementation. Consequently, Newton iterations may stall if an iterate happens to become too large.